OpenStudy (anonymous):

Barney has 17 2/3 feet of lumber that he is going to cut into 4 equal pieces for a border on his garden. Which of the following is true? Each piece of lumber is feet long. a Each piece of lumber is 1 5/12 feet long. b Each piece of lumber is 13 5/12 feet long. c Each piece of lumber is 4 5/12 feet long. dEach piece of lumber is 13 1/2 feet long.

OpenStudy (anonymous):

@destiny12345

OpenStudy (anonymous):

@alicia25

OpenStudy (anonymous):

c is the correct answer 17 2 ÷ 4 = 17·3 + 2 ÷ 4 = 53 ÷ 4 = 53 × 1 = 53·1 = 53 5 + 4·12 = 4 5 3 3 1 3 1 3 4 3·4 12 12 12

OpenStudy (anonymous):

tell if im wrong

OpenStudy (anonymous):

your right

OpenStudy (anonymous):

YYAAYY

OpenStudy (anonymous):

you are good at math @destiny12345 i think your very smart

OpenStudy (anonymous):

im not very good at it but ill take the compliment so thank you

OpenStudy (anonymous):

you are